OpenStudy (firejay5):

Solve the absolute value. There should be 2 answers, if not 1.: 4|3t - 8| = 16t

OpenStudy (anonymous):

This is the same concept. Start by dividing by 4. |3t - 8| = 4t. 3t - 8 = -4t or 3t - 8 = 4t 7t = 8 or t = -8

OpenStudy (firejay5):

I need both answers for them

OpenStudy (anonymous):

t = 8/7 and t = -8

OpenStudy (firejay5):

thank you click on link for other math problem: http://openstudy.com/study#/updates/503d635ee4b0c29115be3462